OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

the second to last one...y-(-5) = 2(x-1) y+5 = 2x-2 y = 2x - 7